A shoe store developed the following estimated regression equation relating sales to inventory investment and advertising expenditures where 1inventory investment ($1000s) = advertising expenditures ($1000s) y sales ($1000s) The data used to develop the model came from a survey of 10 stores; for those data, SST 16,000 and SSR a. Compute SSE, MSE, and MSR (to 2 decimals, if necessary) 12,000 SSE MSE MSR b. Use an F test and α .05 level of significance to determine whether there is a relationship among the variables. Compute F test statistic (to 2 decimals). Use F table. The p-value is| . Select your answer- What is your conclusion? There -Select your answer- a significant relationship among the variables.
Answer #1

a) SSE = SST - SSR = 16000 - 12000 = 4000

MSE = SSE/dfE = 4000/(10 - 3) = 4000/7 = 571.43

MSR = SSR/dfR = 12000/(3 - 1) = 12000/2 = 6000

b) F test statistic = MSR/MSE = 6000/571.43 = 10.50

p - Value is 0.0078 so it is between 0.005 and 0.010

There is a significant relationship.